# Vlaser: Vision-Language-Action Model with Synergistic Embodied Reasoning

### v1

@@ 2.2 Vlaser Data Engine
General and Spatial Reasoning Data The Vlaser dataset integrates 1.2 million question-answer pairs dedicated to general Robotic Visual Question Answering (RoboVQA) tasks, along with an additional 500k data items specifically designed to enhance spatial intelligence. This comprehensive data composition substantially strengthens the model’s capabilities in general state perception and 3D spatial reasoning. For the RoboVQA component, data is aggregated from multiple established sources, including RoboVQA (Sermanet et al., 2024), Robo2VLM (Chen et al., 2025b), RoboPoint (Yuan et al., 2024), RefSpatial (Zhou et al., 2025a), OWMM-Agent (Chen et al., 2025a), among others. To support spatial understanding and reasoning, we incorporate datasets such as SPAR (Zhang et al., 2025), SpaceR-151k (Ouyang et al., 2025), and VILASR (Wu et al., 2025). Furthermore, we augment these with 100k manually annotated spatial understanding samples generated from publicly available 3D scene datasets—including ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), CA-1M (Lazarow et al., 2025), and ARKitScenes (Baruch et al., 2021). The integration of these diverse and high-quality data sources effectively enhances the model’s spatial awareness and supports more robust performance in complex embodied reasoning tasks.
@@
In-Domain Data for downstream VLAs To facilitate the end-to-end policy learning for Vision-Language Action Models (VLAs), we further generate 2 million in-domain multimodal question-answer pairs tailored for VLM pretraining. These data are specifically designed to align with the embodied reasoning context and enhance the model’s ability to perceive, reason, and plan in interactive environments. The in-domain data is sourced from simulation platforms SimplerEnv (Li et al., 2024c). Within SimplerEnv, data is generated for two distinct robotic embodiments: the Google Robot (Brohan et al., 2023b; a; O’Neill et al., 2024) and the WidowX Robot (Walke et al., 2023a), ensuring broad morphological and kinematic coverage. The question-answer pairs encompass the specialized categories including embodied grounding, spatial intelligence, planning and general VQA for robot states as described above. The detailed methodology for constructing each of the in-domain data in simulation is described in Appendix A.2.
@@ 2.3 Training Recipe
Vision-Language-Action Finetuning For robot policy learning, we optimize the model using an action expert trained on robot-specific datasets. Vlaser integrates a flow-matching-based action expert to predict a sequence of future actions from a single-frame observation. Specifically, denote the action chunk , where represents the action in the current timestep and represents the action horizon. Meanwhile, we encode each noisy action with an action encoder (i.e., an MLP projector) as a single token for the action expert. For the action chunk, is the corresponding noisy action chunk, and we train the network to output to match the denoising vector field , where indicates the observations (e.g., image camera and robot state) at action timestep , represents the network and represents the flow matching timesteps. Therefore, the VLA optimization loss is as follows,
We sample the action chunks from the robot episodes and flow-matching timesteps to optimize the network. At inference, we generate actions by integrating the learned vector field from to , starting with random noise , as follows,
where is the integration step size. In our experiments, we set as 4, and as 10. We aim to identify the most effective VLMs for downstream VLA fine-tuning and bridge the gap between foundational VLMs and their performance in downstream VLA tasks, thus shedding light on the future construction of embodied VLMs. Currently, the SimplerEnv benchmark, including Bridge (Walke et al., 2023b) and Google Robot (Jang et al., 2022; Brohan et al., 2023b) datasets, provides numerous training episodes (Over 5M images) and corresponding Real-to-Sim benchmarks. We thus majorly analyze the most effective data stream for VLA finetuning based on SimplerEnv.
@@ 3.1 Performance on Embodied Reasoning Capability
3.2 Performance on downstream Close-Loop Robot Tasks
Finetuning Datasets We conduct extensive experiments on SimplerENV to evaluate the performance of Vlaser and Vlaser data engine on closed-loop robotic manipulation tasks. SimplerENV is an open-source suite of purpose-built simulated environments with nearly 150K episodes for evaluating real-world robot manipulation policies in a scalable, reproducible way. It targets the key real-to-sim gaps – control and vision so that simulated performance reliably tracks real-robot outcomes. Across Google Robot and WidowX/BridgeData V2 setups, SimplerEnv reports strong real-vs-sim correlations and faithfully reflects behavior under distribution shifts, enabling fast, comparable policy assessment without full digital twins. As a result, SimplerENV has been widely adopted for evaluating VLA models and has proven to reliably reflect the performance of the models on the real robot platform.
@@ 3.2 Performance on downstream Close-Loop Robot Tasks
| Vlaser (2B) | 60.8% | 35.4% | 56.7% | 20.0% | 43.2% |
| Vlaser-QA (2B) | 55.8% | 83.3% | 77.9% | 33.3% | 64.6% |
@@
Baselines As mentioned before, we integrate Vlaser with an action expert utilizing flow matching for low-level control fine-tuning and evaluating the model on the WidowX and Google Robot datasets from SimplerEnv. We conduct experiments using InternVL3-2B, Vlaser, and the models based on Vlaser data engine for the in-domain robot data. Commonly, embodied reasoning can be marginally categorized into embodied QA (including planning), embodied grounding, and embodied spatial intelligence. We thus construct corresponding embodied VLMs based on the corresponding data stream: Vlaser-QA, Vlaser-Grounding, Vlaser-Spatial. Notably, all series of Vlaser models are based on the same architecture with 2B size.
The full experimental results are presented in Table 2 and Table 3. We observe the proposed vision-language-action architecture based on InternVL3-2B is capable of low-level robot control and achieves comparable performance to many previous approaches, though we utilize a 2B backbone.
@@
Particularly, we did not observe any clean improvement when we used the Vlaser-2B as our initial backbone on both WidowX and Google Robot, while we achieved significant performance promotion with the Vlaser-QA, although the architecture and model size are the same between the two models. This observation illustrates the effectiveness of our Vlaser data engine, and meanwhile identifies that there is no positive correlation between common embodied reasoning benchmarks and the performance of closed-loop control of the lower level for the specific embodiment of the robot. We reckon it is the domain shift between the internet data and the corresponding robot embodiment (e.g., WidowX or Google Robot), and we find that the enhanced abilities in the same observation domain effectively facilitate the closed-loop success rate. Therefore, it is urgent to shrink the domain gap between the foundational models and real-world robot embodiment for closed-loop task completion.
In addition, we conducted a simple ablation study on the Google Robot Tasks to evaluate the effectiveness of different data annotations. The experimental results indicate that incorporating all types of data leads to significant improvements, achieving performance comparable to the baseline. We attribute these improvements primarily to the reduction of the vision observation domain shift.
@@ 4 Related Work
Vision-Language Model for Embodied Reasoning Enhancing the embodied reasoning capabilities of current state-of-the-art Vision-Language Models (VLMs) has emerged as a critical research direction. These capabilities encompass a range of competencies, including grounding (Yuan et al., 2024; Deitke et al., 2025; Cheng et al., 2025) which identifies affordances that enable embodied agents to perform manipulations, spatial intelligence (Yang et al., 2025b; d), such as object counting and spatial relationship understanding, as well as task planning (Chen et al., 2023; Qiu et al., 2024), which involves assessing the current state and determining subsequent actions to be executed. Gemini Robotics-ER (Team et al., 2025b) integrates embodied reasoning into its core visual-language model (VLM), demonstrating strong generalization across a variety of tasks such as 3D scene perception, visual pointing, state estimation, and affordance prediction. In parallel, a number of data-driven methodologies have emerged to support such reasoning capabilities. For instance, Cosmos-Reason1 (NVIDIA et al., 2025a), VeBrain (Luo et al., 2025), MolmoAct (Lee et al., 2025), and EmbodiedOneVision (Qu et al., 2025a) each contribute curated datasets specifically designed for embodied reasoning tasks, emphasizing aspects such as multi-modal instruction following and action-aware visual-language alignment. Further advancing this direction, several frameworks – including the RoboBrain series (Ji et al., 2025; Team et al., 2025a), Embodied R1 (Yuan et al., 2025), and Robix (Fang et al., 2025) incorporate Reinforcement Fine-Tuning (RFT) and synthesize spatio-temporal reasoning datasets enriched with structured thought traces. These approaches aim to enhance models’ capacity for causal reasoning and long-horizon task decomposition. Distinguished from these prior efforts, our work not only achieves competitive, and in some cases superior performance on established embodied reasoning benchmarks, but also provides an in-depth analysis of the synergistic relationship between pre-trained VLMs and downstream Vision-Language Action Models (VLAs), offering insights that bridge model capabilities and real-world deployment.
@@
Vision-Language-Action models. Developing a generalist model remains a central challenge in robotics. Inspired by the strong generalization abilities of vision-language models (VLMs) (OpenAI, 2023; Chen et al., 2024; Bai et al., 2025; Team et al., 2023) trained on large-scale internet data, researchers have proposed vision-language-action (VLA) models, which have demonstrated promising performance (Brohan et al., 2023a; Kim et al., 2024; Qu et al., 2025b; Hou et al., 2025). Compared to traditional robot policies, VLA models are pretrained on large-scale robotics datasets and exhibit improved generalization across object categories and visual observations. Building on recent progress, researchers have incorporated techniques such as diffusion (Ho et al., 2020; Rombach et al., 2022; Peebles & Xie, 2023), flow matching (Lipman et al., 2023b), and mixture-of-experts (MoE) (Shazeer et al., 2017) into VLA models, and have adopted larger, more capable VLMs as their backbones. These advances have enabled VLA models to tackle a wider range of complex real-world manipulation tasks. Nevertheless, current VLA models remain limited in their generalization. In particular, they have not yet reached the level of general reasoning exhibited by VLMs, such as decompose a complex task into manageable sub-steps and complete the task in a zero-shot manner. Efforts have been made to enhance specific reasoning abilities in embodied scenarios (Team et al., 2025b; Ji et al., 2025; NVIDIA et al., 2025a). In parallel, several studies (Intelligence et al., 2025; Driess et al., 2025; Zhou et al., 2025b) explore unified training frameworks for VLMs and VLAs to leverage the reasoning capacity of VLMs. However, the relationship between high-level multimodal reasoning and low-level control performance remains largely unexplored. It is still unclear which specific multimodal abilities such as spatial understanding, grounding, or planning and which types of training data most effectively enhance the control capabilities of a VLA model. In this work, we take an initial step toward analyzing this relationship by a systematic evaluation, and also propose the latest foundational model with strong embodied multimodal understanding and action prediction.

### v2

@@ 2.2 Vlaser Data Engine
General and Spatial Reasoning Data The Vlaser dataset integrates 1.2 million question-answer pairs dedicated to general Robotic Visual Question Answering (RoboVQA) tasks, along with an additional 500k data items specifically designed to enhance spatial intelligence. This comprehensive data composition substantially strengthens the model’s capabilities in general state perception and 3D spatial reasoning. For the RoboVQA component, data is aggregated from multiple established sources, including RoboVQA (Sermanet et al., 2024), Robo2VLM (Chen et al., 2025b), RoboPoint (Yuan et al., 2024), RefSpatial (Zhou et al., 2025a), OWMM-Agent (Chen et al., 2025a), among others. To support spatial understanding and reasoning, we incorporate open-source datasets such as SPAR (Zhang et al., 2025), SpaceR-151k (Ouyang et al., 2025), and VILASR (Wu et al., 2025). Furthermore, we augment these with 100k manually annotated spatial understanding samples generated from publicly available 3D scene datasets—including ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), CA-1M (Lazarow et al., 2025), and ARKitScenes (Baruch et al., 2021). The integration of these diverse and high-quality data sources effectively enhances the model’s spatial awareness and supports more robust performance in complex embodied reasoning tasks.
@@
In-Domain Data for downstream VLAs To facilitate the end-to-end policy learning for Vision-Language Action Models (VLAs), we further generate 2 million in-domain multimodal question-answer pairs tailored for VLM pretraining. These data are specifically designed to align with the embodied reasoning context and enhance the model’s ability to perceive, reason, and plan in interactive environments. The in-domain data is sourced from simulation platforms SimplerEnv (Li et al., 2024d) and RoboTwin (Chen et al., 2025c). Within SimplerEnv, data is generated for two distinct robotic embodiments: the Google Robot (Brohan et al., 2023b; a; O’Neill et al., 2024) and the WidowX Robot (Walke et al., 2023a), and within RoboTwin, data is generated from dual-arm Aloha-AgileX Robot. The question-answer pairs encompass the specialized categories including embodied grounding, spatial intelligence, planning and general VQA for robot states as described above. The detailed methodology for constructing and filtering each of the in-domain data in simulation is described in Appendix A.2.
@@ 2.3 Training Recipe
Vision-Language-Action Finetuning For robot policy learning, we optimize the model using an additional incorporated action expert module trained on robot-specific datasets. The action expert is analogous to a mixture of experts(MoE) (Shazeer et al., 2017b; Du et al., 2022; Zhou et al., 2024) architecture with two mixture elements, while the original part of parameter is used for image and text inputs, and the additionally separate set of weights for the robotics-specific (action and state) tokens inputs and outputs are referred as the action expert. Vlaser integrates a flow-matching-based action expert to predict a sequence of future actions from a single-frame observation. Specifically, we denote the action chunk , where represents the action in the current timestep and represents the action horizon. indicates the observations (image with views, language and robot state ) at action timestep . , and are encoded via corresponding encoders and then projected via a linear projection layer into the same embedding space. represents the action expert network and represents the flow matching timesteps. For the action chunk, is the corresponding noisy action chunk, and we train the network to output to match the denoising vector field where . Therefore, the VLA optimization loss is as follows,
Formally, following prior flow-matching based VLA works (Black et al., 2024; Zren, 2025), We sample the action chunks from the robot episodes and flow-matching timesteps to optimize the network. At inference, we generate actions by integrating the learned vector field from to , starting with random noise , as follows,
where is the integration step size. In our experiments, we set as 4, and as 0.1( integration steps) at inference time for the improvement of inference efficiency. We aim to identify the most effective VLMs for downstream VLA fine-tuning and bridge the gap between foundational VLMs and their performance in downstream VLA tasks, thus shedding light on the future construction of embodied VLMs. Currently, the SimplerEnv benchmark, including Bridge (Walke et al., 2023b) and Google Robot (Jang et al., 2022; Brohan et al., 2023b) datasets, provides numerous training episodes (Over 5M images) and corresponding Real-to-Sim benchmarks. We thus majorly analyze the most effective data stream for VLA finetuning based on SimplerEnv.
@@ 3.1 Performance on Embodied Reasoning Capability
| Vlaser-OOD (2B) | 60.8% | 35.4% | 56.7% | 20.0% | 43.2% |
| Vlaser-QA (2B) | 55.8% | 83.3% | 77.9% | 33.3% | 62.6% |
| Vlaser-Spatial (2B) | 48.3% | 81.7% | 76.7% | 36.7% | 60.8% |
| Vlaser-Grounding (2B) | 47.5% | 80.8% | 80.0% | 39.6% | 62.0% |
| Vlaser-All (2B) | 52.5% | 87.9% | 76.6% | 43.3% | 65.1% |
3.2 Performance on downstream Close-Loop Robot Tasks
Finetuning Datasets We firstly conduct extensive experiments on SimplerENV (Li et al., 2024d) to evaluate the performance of Vlaser and Vlaser data engine on closed-loop robotic manipulation tasks. SimplerENV is an open-source suite of purpose-built simulated environments with nearly 150K episodes for evaluating real-world robot manipulation policies in a scalable, reproducible way. It targets the key real-to-sim gaps – control and vision so that simulated performance reliably tracks real-robot outcomes. Across Google Robot and WidowX/BridgeData V2 setups, SimplerEnv reports strong real-vs-sim correlations and faithfully reflects behavior under distribution shifts, enabling fast, comparable policy assessment without full digital twins. As a result, SimplerENV has been widely adopted for evaluating VLA models and has proven to reliably reflect the performance of the models on the real robot platform. To further demonstrate the generalizability of our method to other simulation platforms and embodiments, we also conduct experiments on RoboTwin (Mu et al., 2024; Chen et al., 2025c) platforms with Aloha-AgileX as bimanual embodiement. Robotwin is a scalable framework for bimanual manipulation, which integrates scalable training sets and pre-defined tasks as benchmarks for comprehensive robust bimanual manipulation.
@@ 3.2 Performance on downstream Close-Loop Robot Tasks
| Vlaser-All (2B) | 91.0% | 85.4% | 52.1% | 76.2% | 80.5% | 77.7% | 18.8% | 59.0% |
| Simulation Task | RDT-1B (Liu et al., 2024) | InternVL3-2B | Vlaser-OOD(2B) | Vlaser-QA (2B) | Vlaser-Spatial (2B) | Vlaser-Grounding (2B) | Vlaser-All (2B) |
| Beat block hammer | 28% | 12% | 20% | 18% | 20% | 32% | 40% |
| Click bell | 46% | 78% | 94% | 48% | 98% | 86% | 92% |
| Handover mic | 92% | 74% | 56% | 84% | 60% | 52% | 84% |
| Move can pot | 44% | 40% | 42% | 66% | 56% | 50% | 46% |
| Move pillbottles pad | 10% | 62% | 70% | 66% | 78% | 68% | 72% |
| Move playingcard away | 20% | 64% | 52% | 58% | 68% | 84% | 74% |
| Pick diverse bottles | 2% | 30% | 44% | 36% | 24% | 34% | 38% |
| Place burger fries | 42% | 36% | 46% | 88% | 82% | 46% | 42% |
| Place container plate | 82% | 72% | 70% | 84% | 78% | 82% | 84% |
| Place phone stand | 8% | 42% | 40% | 38% | 36% | 48% | 50% |
| Place mouse pad | 2% | 68% | 30% | 44% | 38% | 48% | 92% |
| Shake bottle | 66% | 92% | 90% | 98% | 96% | 98% | 96% |
| Avg. | 36.8% | 55.8% | 54.5% | 60.7% | 61.2% | 60.7% | 67.5% |
Baselines Alongside comparisons with other commonly used VLA models (Black et al., 2024; Kim et al., 2024; Liu et al., 2025), we conduct a clear self-comparable ablation study to evaluate the individual contributions of different in-domain data sources. Specifically, InternVL3-2B denotes the base InternVL model, while Vlaser-OOD refers to the Vlaser-2B model fine-tuned solely on Vlaser-6M out-of-domain (OOD) data specific for the embodied reasoning benchmarks in Sec. 3.1, without any in-domain data in Vlaser-6M dataset. Regarding the in-domain datas derived directly from the simulation platform, we categorize them into three types: embodied QA (including planning), embodied spatial intelligence, and embodied grounding. The corresponding fine-tuned Vlaser-2B models are denoted as Vlaser-QA, Vlaser-Spatial, and Vlaser-Grounding, respectively. Additionally, we combine all the three in-domain data types to fine-tune a model referred to as Vlaser-All.
@@
The full experimental results are presented in Table 2, Table 3 and Table 4. Notably, no clear improvement was observed when using Vlaser-OOD-2B as the initial backbone across all three benchmarks. The task success rates of Vlaser-OOD and the baseline InternVL3-2B remain close. Conversely, all models fine-tuned with in-domain data—Vlaser-QA, Vlaser-Spatial, Vlaser-Grounding, and Vlaser-All—exhibit significant performance gains, even though the model architecture and size remain unchanged. This observation illustrates the effectiveness of our Vlaser data engine, and meanwhile identifies that there is no positive correlation between common embodied reasoning benchmarks and the performance of closed-loop control of the lower level for the specific embodiment of the robot. We reckon it is the domain shift between the internet data and the corresponding robot embodiment (e.g., WidowX or Google Robot), and we find that the enhanced abilities in the same observation domain effectively facilitate the closed-loop success rate. Therefore, it is urgent to shrink the domain gap between the foundational models and real-world robot embodiment for closed-loop task completion.
Regarding the three types of in-domain data annotations, we experimentally find that incorporating any of them leads to significant performance gains over the baseline. We attribute these improvements primarily to the mitigation of visual observation domain shift. Furthermore, integrating all three types of in-domain data results in further performance enhancement. This suggests that pretraining with diverse in-domain multimodal data, spanning general QA, grounding, and spatial intelligence, could best facilitates transfer learning for VLA policy learning and leads to improved task success rates.
3.3 Ablation Studies
In this section, we adopt ablation studies regarding three key hyperparameters for VLA end-to-end training, i.e.,, the predicted action length , the execute action length sizes as well as flow-matching sampling steps . By default we use as 4, with as 4, and as 10.
| Model | Predict Length | Execute Length | Sample Steps | Carrot on plate | Put eggplant in basket | Spoon on the towel | Stack cube | Avg |
| InternVL-2B | 4 | 4 | 10 | 42.9% | 57.1% | 55.8% | 11.3% | 41.8% |
| 4 | 2 | 10 | 22.9% | 18.3% | 40.8% | 2.9% | 21.2% | |
| 2 | 2 | 10 | 34.6% | 22.9% | 54.2% | 2.9% | 28.7% | |
| 4 | 4 | 20 | 38.8% | 54.2% | 51.3% | 8.3% | 38.2% | |
| Vlaser-OOD(2B) | 4 | 4 | 10 | 60.8% | 35.4% | 56.7% | 20.0% | 43.2% |
| 4 | 2 | 10 | 50.0% | 21.7% | 30.0% | 12.1% | 28.5% | |
| 2 | 2 | 10 | 62.5% | 19.2% | 49.2% | 21.3% | 38.1% | |
| 4 | 4 | 20 | 57.5% | 29.2% | 54.6% | 17.1% | 39.6% | |
| Vlaser-QA(2B) | 4 | 4 | 10 | 55.8% | 83.3% | 77.9% | 33.3% | 62.6% |
| 4 | 2 | 10 | 44.2% | 64.2% | 59.6% | 36.3% | 51.1% | |
| 2 | 2 | 10 | 47.5% | 66.3% | 67.1% | 36.3% | 54.3% | |
| 4 | 4 | 20 | 56.3% | 85.0% | 76.7% | 35.0% | 63.3% |
The results are shown in Table 5. It is clear to see that, no matter in which group of hyperparameter settings, the performance based on Vlaser-OOD shows slight improvement compared with the performance based on InternVL3-2B. Besides, there is a significant gains while using the model fine-tuned with in-domain data Vlaser-QA. This conclusion is as same as the results in 3.2, which demonstrates great robustness of our method.
@@ 4 Related Work
Vision-Language Model for Embodied Reasoning Enhancing the embodied reasoning capabilities of current state-of-the-art Vision-Language Models (VLMs) has emerged as a critical research direction. These capabilities encompass a range of competencies, including grounding (Yuan et al., 2024; Deitke et al., 2025; Cheng et al., 2025) which identifies affordances that enable embodied agents to perform manipulations, spatial intelligence (Yang et al., 2025b; d), such as object counting and spatial relationship understanding, as well as task planning (Chen et al., 2023; Qiu et al., 2024), which involves assessing the current state and determining subsequent actions to be executed. Gemini Robotics-ER (Team et al., 2025b) integrates embodied reasoning into its core visual-language model (VLM), in parallel, a number of data-driven methodologies have emerged to support such reasoning capabilities. For instance, Cosmos-Reason1 (NVIDIA et al., 2025a), VeBrain (Luo et al., 2025), MolmoAct (Lee et al., 2025), and EmbodiedOneVision (Qu et al., 2025a) each contribute curated datasets specifically designed for embodied reasoning tasks, emphasizing aspects such as multi-modal instruction following and action-aware visual-language alignment. Furthermore, several frameworks – including the RoboBrain series (Ji et al., 2025; Team et al., 2025a), Embodied R1 (Yuan et al., 2025), and Robix (Fang et al., 2025) incorporate Reinforcement Fine-Tuning (RFT) and synthesize spatio-temporal reasoning datasets enriched with structured thought traces. These approaches aim to enhance models’ capacity for causal reasoning and long-horizon task decomposition. Distinguished from these prior efforts, our work not only achieves competitive, and in some cases superior performance on established embodied reasoning benchmarks, but also provides an in-depth analysis of the synergistic relationship between pre-trained VLMs and downstream Vision-Language Action Models (VLAs), offering insights that bridge model capabilities and real-world deployment.
@@
Vision-Language-Action models. Developing a generalist model remains a central challenge in robotics. Inspired by the strong generalization abilities of vision-language models (VLMs) (OpenAI, 2023; Chen et al., 2024; Bai et al., 2025; Team et al., 2023) trained on large-scale internet data, researchers have proposed vision-language-action (VLA) models, which have demonstrated promising performance (Brohan et al., 2023a; Kim et al., 2024; Qu et al., 2025b; Hou et al., 2025). Compared to traditional robot policies, VLA models are pretrained on large-scale robotics datasets and exhibit improved generalization across object categories and visual observations. Building on recent progress, researchers have incorporated techniques such as diffusion (Ho et al., 2020; Rombach et al., 2022; Peebles & Xie, 2023), flow matching (Lipman et al., 2023b), and mixture-of-experts (MoE) (Shazeer et al., 2017a) into VLA models, and have adopted larger, more capable VLMs as their backbones. These advances have enabled VLA models to tackle a wider range of complex real-world manipulation tasks. Efforts have been made to enhance specific reasoning abilities in embodied scenarios (Team et al., 2025b; Ji et al., 2025; NVIDIA et al., 2025a). In parallel, several studies (Intelligence et al., 2025; Driess et al., 2025; Zhou et al., 2025b) explore unified training frameworks for VLMs and VLAs to leverage the reasoning capacity of VLMs. However, the relationship between high-level multimodal reasoning and low-level control performance remains largely unexplored. It is still unclear which specific multimodal abilities such as spatial understanding, grounding, or planning and which types of training data most effectively enhance the control capabilities of a VLA model. In this work, we take an initial step toward analyzing this relationship by a systematic evaluation, and also propose the latest foundational model with strong embodied multimodal understanding and action prediction.
5 Limitations
One limitation of this work is the absence of real-robot experiments. While the experiments conducted in this study have all been performed well in simulation, in future work we plan to conduct more experiments with real robots to further evaluate and refine the proposed methods.